# SMART: Scalable Multi-agent Real-time Simulation via Next-token Prediction

### v1

@@ Abstract
Data-driven autonomous driving motion generation tasks are frequently impacted by the limitations of dataset size and the domain gap between datasets, which precludes their extensive application in real-world scenarios. To address this issue, we introduce SMART, a novel autonomous driving motion generation paradigm that models vectorized map and agent trajectory data into discrete sequence tokens. These tokens are then processed through a decoder-only transformer architecture to train for the next token prediction task across spatial-temporal series. This GPT-style method allows the model to learn the motion distribution in real driving scenarios. SMART achieves state-of-the-art performance across most of the metrics on the generative Sim Agents challenge, ranking 1st on the leaderboards of Waymo Open Motion Dataset (WOMD), demonstrating remarkable inference speed. Moreover, SMART represents the generative model in the autonomous driving motion domain, exhibiting zero-shot generalization capabilities: Using only the NuPlan dataset for training and WOMD for validation, SMART achieved a competitive score of 0.71 on the Sim Agents challenge. Lastly, we have collected over 1 billion motion tokens from multiple datasets, validating the model’s scalability. These results suggest that SMART has initially emulated two important properties: scalability and zero-shot generalization, and preliminarily meets the needs of large-scale real-time simulation applications. We have released all the code to promote the exploration of models for motion generation in the autonomous driving field.
@@ 3.2 Model Architecture
where $M_{i}$ denotes the neighbor set of the road tokens. To incorporate spatial awareness for map encoding, we generate the $j^{th}$ key/value vector from the concatenation of $r_{j}$ and the relative positional embedding $\text{RPE}_{ij}$cui2023gorela .
@@
$e_{i^{\prime}}=\text{MHSA}\left(q(e_{i}^{t}),k(e_{i}^{t-\tau}+\text{RPE}_{i}^{t,t-\tau}),v(e_{i}^{t-\tau},\text{RPE}_{i}^{t,t-\tau})\right),\quad 0<\tau<t$ (2a)
@@ 3.3 Spatial-temporal next token prediction
As shown in Figure 2(b), the road vector NTP task targets RoadNet to learn the spatial structure of road vector inputs. Unlike agent motions, road vectors form a graph rather than a sequence, making it challenging to apply next token prediction tasks directly. To address this issue, we extract the original topological information of roads and model the road vector tokens with sequential relationships based on their predecessor-successor connections. As depicted in Figure 2(b), in the pre-training NTP task, the subsequent road vector token is predicted using the preceding road token based on the road topology. This approach requires RoadNet to understand the connectivity and continuity among unordered road vectors.
@@
$loss(\theta)=-\sum_{t=1}^{T}\sum_{i=1}^{V_{a}}\left(a_{i}^{t+1}==a_{i^{gt}}^{t+1})log(p(a_{i}^{t+1}|e_{i}^{1:t,r_{j}}),\right.$ (3)
@@ 4 Experiments
To validate the generalizability and scalability of the SMART model, we conducted extensive experiments and trained models across various scales. On the official WOMD Sim Agents Challenge (WOSAC), we employed the SMART 8 Million parameters (8M) model, which was exclusively trained on the WOMD dataset. Concurrently, the SMART 8M model was also utilized for generalization experiments and ablation studies. In the scale law experiments, we integrated additional datasets and trained on models of multiple scales. For all experiments, the testing datasets employed the split validation dataset from WOMD. Detailed hyperparameters for the SMART architecture can be found in Section A.1. In the following sections, Section 4.1 presents the results of rollouts generated by SMART on the WOSAC benchmark montali2024waymo . Evaluations of SMART’s generalizability and scalability are detailed in Sections 4.2 and 4.3, respectively. Finally, an ablation analysis of our design methods is conducted in Section 4.4.
@@ 4.1 Comparison for motion generation task
We compare proposed SMART with existing motion generation approaches including diffusion modelsguo2023scenedm , continuous distribution regression models wang2023multiverse ; shi2023mtr++ , and next token autoregressive modelphilion2023trajeglish . Because the Sim Agents challenge metrics were changed twice, to compare it more broadly with the previous methodology, we test the performance of our model using both the WOMD Sim Agents 2023 and 2024 Benchmarkmontali2024waymo . As shown in Table 1 and Table 2, SMART achieves not only the best Realism Meta metric but also a high prediction precision. Most of the improvement is because SMART models interaction between agent-agent and map-agent significantly better than prior work. For further detailed comparisons, please refer to A.2.
@@ 4.2 Generalization
Figure 3: Qualitative results of closed-loop planning for two representative scenarios from the test set. Each scenario (every row) lasts 8 seconds and we take 4 snapshots with a 2-second interval. All the agents in the scenario are controlled by SMART. The first row depicts a parking lot area. The red vehicle in the picture effectively completed a detour around a stationary vehicle ahead in the parking lot. The second row shows a scene of a large curvature U-turn in a ramp zone, where the traffic flow in the right lane of the ramp has completed the behavior of ramp exit under the control of SMART. It is recommended to refer to supplementary materials for more videos
@@ 4.3 Scalability
Figure 4: Due to limitations in dataset size, we trained models at multiple scales ranging from 1M to 96M on a total of 1 billion tokens. (a) Training loss of different models (b) Axes are all on a logarithmic scale. The power-law scaling law can be expressed as a solid line. Exponents $\beta=-0.157$ suggest a smooth decline in test loss $L$ when scaling up SMART models.
@@ 4.4 Ablation
In this study, we aim to verify the effectiveness of each component of SMART. Results are reported in Table 4. The initial model, denoted as $M1$, is constructed on the architecture depicted in Sec.3.2, employing solely agent tokenization. The introduction of the road vector tokenization in $M2$, which tokenized the road vector states into discrete tokens, results in marked improvements over $M1$ in the generalization capability. Comparing models $M1$ and $M2$ reveals that when trained solely on the WOMD dataset, the tokenization of road vectors results in a certain reduction in overall metrics. We speculate that discretized map tokens may lose some fine-grained geometric information about roads. $M3$ incorporates noised agent motion tokenization, designed to address cumulative errors and distributional shifts during inference. This modification leads to enhancements in both the interaction metric and the map-based metric.
Table 4: Ablation study on each component of SMART. Experimental results are based on the WOMD validation set. "RVT" indicates road vector tokenization, "NAT" indicates noised agent tokenization, "NRVT" indicates noised road vector tokenization
@@
$M_{4}$ $\surd$ $\surd$ $\surd$ 0.453 0.803 0.851 0.416 0.785 0.797
@@ A.1 Implementation and Simulation Inference
SMART 1M SMART 8M SMART 36M SMART 96M
RoadNet Number of self attention layers 1 1 1 2
Road token embeddings 32 128 256 512
@@
Road token attention radius 40 40 40 40
MotionNet Number of temporal attention layers 1 3 3 4
Number of agent-agent attention layers 1 3 3 4
Number of map-agent attention layers 1 3 3 4
Number of attention head 4 8 8 8
@@
Feature dimension of Agent token embeddings 32 128 256 512
Size of motion token vocabulary 512 512 512 2048
SMART Total parameters 1.1M 8.1M 36.7M 96.3M
@@ A.2 Detailed comparison in the WOSAC leaderboard
SMART 8M 0.363 0.296 0.423 0.564 0.376 0.963 0.832 0.659 0.936 1.54
Here’s the revised version of your text with improved precision and grammar:
@@
In the benchmark comparisons presented in Table 6, the SMART 8M method, developed by our team, demonstrates superior performance across multiple metrics, particularly excelling in interactive and safety-related indicators. Notably, SMART 8M achieved the highest scores in angular acceleration, distance to nearest object, collision avoidance, and off-road metrics, underscoring its effectiveness in complex driving scenarios. These results highlight the robustness of SMART 8M in ensuring safety and reliability, indicating its advanced capability in managing dynamic and potentially hazardous traffic conditions more effectively than other evaluated methods. This performance also suggests the potential of the SMART model to be applied to planning tasks.
@@ A.3 Additional ablation studies
Table 7: Comparison of different tokenizer. Experimental results are based on the SMART 8M
@@
SMART 1M 0.417 0.761 0.828 8hours 10.30ms
SMART 8M 0.453 0.803 0.851 23hours 10.79ms
SMART 36M 0.456 0.810 0.861 3days 19.73ms
SMART 96M 0.457 0.819 0.872 1week 46.58ms
The results in Table 9 highlight the performance of SMART models with different parameter scales across various metrics. As the model scale increases from SMART 1M to SMART 96M, there is a significant improvement in both the interactive metrics and the map-based metrics. This indicates that larger models are better at capturing interactions and understanding map-based context, leading to enhanced performance in these areas. However, the kinematic metrics show minimal variation. Additionally, the training time and average inference time increase substantially with larger models, reflecting the trade-off between model performance and computational cost. Validation is conducted every 50,000 train steps. The model is considered to have converged if there is no significant loss reduction or metric improvement after five consecutive validations. The training and inference time is measured on 32 NVIDIA TESLA V100 GPUs.

### v3

@@ Abstract
Data-driven autonomous driving motion generation tasks are frequently impacted by the limitations of dataset size and the domain gap between datasets, which precludes their extensive application in real-world scenarios. To address this issue, we introduce SMART, a novel autonomous driving motion generation paradigm that models vectorized map and agent trajectory data into discrete sequence tokens. These tokens are then processed through a decoder-only transformer architecture to train for the next token prediction task across spatial-temporal series. This GPT-style method allows the model to learn the motion distribution in real driving scenarios. SMART achieves state-of-the-art performance across most of the metrics on the generative Sim Agents challenge, ranking 1st on the leaderboards of Waymo Open Motion Dataset (WOMD), demonstrating remarkable inference speed. Moreover, SMART represents the generative model in the autonomous driving motion domain, exhibiting zero-shot generalization capabilities: Using only the NuPlan dataset for training and WOMD for validation, SMART achieved a competitive score of 0.72 on the Sim Agents challenge. Lastly, we have collected over 1 billion motion tokens from multiple datasets, validating the model’s scalability. These results suggest that SMART has initially emulated two important properties: scalability and zero-shot generalization, and preliminarily meets the needs of large-scale real-time simulation applications. We have released all the code to promote the exploration of models for motion generation in the autonomous driving field. The source code is available at https://github.com/rainmaker22/SMART.
@@ 3.2 Model Architecture
where $R_{i}$ denotes the neighbor set of the road tokens. To incorporate spatial awareness for map encoding, we generate the $j^{th}$ key/value vector from the concatenation of $r_{j}$ and the relative positional embedding $\text{RPE}_{ij}$cui2023gorela .
@@
$\begin{split}e_{i^{\prime}}=\text{MHSA}\left(q(e_{i}^{t}),k(e_{i}^{t-\tau},\text{RPE}_{i}^{t,t-\tau}),v(e_{i}^{t-\tau},\text{RPE}_{i}^{t,t-\tau})\right),\quad 0<\tau<t\end{split}$ (2a)
@@ 3.3 Spatial-temporal next token prediction
As shown in Figure 2(b), the road vector NTP task targets RoadNet to learn the spatial structure of road vector inputs. Unlike agent motions, road vectors form a graph rather than a sequence, making it challenging to apply next token prediction tasks directly. To address this issue, we extract the original topological information of roads and model the road vector tokens with sequential relationships based on their predecessor-successor connections. As depicted in Figure 2(b), in the pre-training NTP task, the subsequent road vector token is predicted using the preceding road token based on the road topology. This approach requires RoadNet to understand the connectivity and continuity among unordered road vectors. The loss function for a single tokenized road vector sequence is defined as:
$\text{loss}(\gamma)=-\sum_{j=1}^{J}\sum_{i=1}^{V_{r}}(\text{r}_{i}^{j+1}==\text{r}_{i^{gt}}^{j+1})\log(p_{\gamma}(\text{r}_{i}^{j+1}|\text{r}_{i}^{1:j}))$ (3)
where $p_{\gamma}(\text{r}_{i}^{j+1}|\text{r}_{i}^{1:j})$ denotes the categorical distribution predicted by the RoadNet parameterized by $\gamma$, $J$ represents a complete polyline that has not yet been split into road vector tokens, $\text{r}^{1:j}$ Representing the road token embedding of the predecessor, and $\text{r}_{i}^{j+1}$ is the next predicted road vector token. This loss function ensures that RoadNet learns to predict the correct next road vector token given the preceding tokens, thereby capturing the spatial continuity and connectivity within the road network.
@@
$loss(\theta)=-\sum_{t=1}^{T}\sum_{i=1}^{V_{a}}(a_{i}^{t+1}==a_{i^{gt}}^{t+1})log(p_{\theta}(a_{i}^{t+1}|e_{i}^{1:t},r_{j}))$ (4)
@@ 4 Experiments
To validate the generalizability and scalability of the SMART model, we conducted extensive experiments and trained models across various scales. On the official WOMD Sim Agents Challenge (WOSAC), we employed the SMART 7 Million parameters (7M) model, which was exclusively trained on the WOMD dataset. Concurrently, the SMART 7M model was also utilized for generalization experiments and ablation studies. In the scale law experiments, we integrated additional datasets and trained on models of multiple scales. For all experiments, the testing datasets employed the split validation dataset from WOMD. Detailed hyperparameters for the SMART architecture can be found in Section A.1. In the following sections, Section 4.1 presents the results of rollouts generated by SMART on the WOSAC benchmark montali2024waymo . Evaluations of SMART’s generalizability and scalability are detailed in Sections 4.2 and 4.3, respectively. Finally, an ablation analysis of our design methods is conducted in Section 4.4.
@@ 4.1 Comparison for motion generation task
We compare proposed SMART with existing motion generation approaches including diffusion modelsguo2023scenedm , continuous distribution regression models wang2023multiverse ; shi2023mtr++ , and next token autoregressive modelphilion2023trajeglish . Because the Sim Agents challenge metrics were changed twice, to compare it more broadly with the previous methodology, we test the performance of our model using both the WOMD Sim Agents 2023 and 2024 Benchmarkmontali2024waymo . As shown in Table 1 and Table 2, SMART achieves not only the best Realism Meta metric but also a high prediction precision. SMART’s modeling approach for maps and motion enables it to learn the behavioral distribution within the data more effectively than prior work. Notably, SMART-zeroshot represents a model trained solely on the NuPlan dataset and directly inferred on the Waymo test set. As shown in Table 2, it achieves performance close to that of MVTE. For further detailed comparisons, please refer to A.2.
@@ 4.2 Generalization
Figure 3: Qualitative results of closed-loop planning for two representative scenarios from the test set. Each scenario (every row) lasts 8 seconds and we take 4 snapshots with a 2-second interval. SMART controls all the agents in the scenario. The first row depicts a parking lot area. The red vehicle in the picture effectively completed a detour around a stationary vehicle ahead in the parking lot. The second row shows a scene of a large curvature U-turn in a ramp zone, where the traffic flow in the right lane of the ramp has completed the behavior of ramp exit under the control of SMART. It is recommended to refer to supplementary materials for more videos
@@ 4.3 Scalability
Figure 4: Due to limitations in dataset size, we trained models at multiple scales ranging from 1M to 101M on a total of 1 billion tokens. (a) Training loss of different models (b) Axes are all on a logarithmic scale. The power-law scaling law can be expressed as a solid line. Exponents $\beta=-0.157$ suggest a smooth decline in test loss $L$ when scaling up SMART models.
@@ 4.4 Ablation
In this study, we aim to verify the effectiveness of each component of SMART. Results are reported in Table 4. The initial model, denoted as $M1$, is constructed on the architecture depicted in Sec.3.2, employing solely agent tokenization. The introduction of the road vector tokenization in $M2$, which tokenized the road vector states into discrete tokens, results in marked improvements over $M1$ in the generalization capability. Comparing models $M1$ and $M2$ reveals that when trained solely on the WOMD dataset, the tokenization of road vectors results in a certain reduction in overall metrics. We speculate that discretized map tokens may lose some fine-grained geometric information about roads. $M4$ incorporates noised agent motion tokenization, designed to address cumulative errors and distributional shifts during inference. This modification leads to enhancements in both the interaction metric and the map-based metric.
Table 4: Ablation study on each component of SMART. Experimental results are based on the WOMD validation set. "RVT" indicates road vector tokenization, "RVNTP" indicates road vector next token prediction, "NAT" indicates noised agent tokenization, "NRVT" indicates noised road vector tokenization
@@
$M_{4}$ $\surd$ $\surd$ $\surd$ 0.437 0.801 0.837 0.411 0.747 0.741
$M_{5}$ $\surd$ $\surd$ $\surd$ 0.453 0.813 0.853 0.413 0.780 0.785
$M_{6}$ $\surd$ $\surd$ $\surd$ $\surd$ 0.453 0.803 0.851 0.416 0.785 0.797
@@ 5 Conclusions
Acknowledgement
We thank the anonymous reviewers, area chairs, and program committee members for their valuable suggestions, which have greatly improved the quality of our work. We also appreciate the thoughtful discussions with Yue Gong and Shuxiang Lu. Authors Wei Wu, Xiaoxin, and Ziyan contributed equally to this work. Wei Wu led the project and provided funding support, while Xiaoxin and Ziyan focused on algorithm design, implementation, model training, and manuscript writing.
@@ A.1 Implementation and Simulation Inference
SMART 1M SMART 7M SMART 26M SMART 101M
RoadNet Number of self attention layers 1 3 1 3
Road token embeddings 64 128 256 512
@@
Road token attention radius 10 10 10 10
MotionNet Number of temporal attention layers 1 6 6 6
Number of agent-agent attention layers 2 6 6 6
Number of map-agent attention layers 2 6 6 6
Number of attention head 8 8 8 8
@@
Feature dimension of Agent token embeddings 64 128 256 512
Size of motion token vocabulary 512 1024 1024 2048
SMART Total parameters 1.0M 7.2M 26.9M 101.0M
@@ A.2 Detailed comparison in the WOSAC leaderboard
SMART 7M 0.363 0.296 0.423 0.564 0.376 0.963 0.832 0.659 0.936 1.749
@@
In the benchmark comparisons presented in Table 6, the SMART 7M method, developed by our team, demonstrates superior performance across multiple metrics, particularly excelling in interactive and safety-related indicators. Notably, SMART 7M achieved the highest scores in angular acceleration, distance to nearest object, collision avoidance, and off-road metrics, underscoring its effectiveness in complex driving scenarios. These results highlight the robustness of SMART 7M in ensuring safety and reliability, indicating its advanced capability in managing dynamic and potentially hazardous traffic conditions more effectively than other evaluated methods. This performance also suggests the potential of the SMART model to be applied to planning tasks.
@@ A.3 Additional ablation studies
Comparison of the scalability and generalization of different architectures
This section presents experiments comparing the architecture proposed in this paper with the MVTE model. The MVTE model [^2] Since MVTE does not have open-source code, we reproduced the results by relying on the MTR model, derived from MTR, represents continuous distribution regression models.
Figure 5: SMART w/o refers to SMART model without the road vector tokenization and noise tricks proposed in this paper. To ensure the fairness of the experiments, all model parameters were adjusted to the 90-100M. Models were trained on various datasets and validated only on the WOMD validation dataset
The experimental results shown in Figure 5 indicate that distribution regression-based models have poor generalization capabilities across different datasets. Models trained with incremental data from other datasets performed worse overall than models trained solely on the WOMD. An interesting phenomenon is that although our private dataset contains more data than the NuPlan dataset, the performance of MVTE trained on it was inferior to that on the NuPlan dataset. This suggests that the distribution regression-based paradigm is likely to cause the model to overfit to a dataset. From the SMART w/o results, it can be seen that the model’s generalization performance is poor, but the effect of incremental data can improve the performance compared to a single training dataset. Based on the above experiments, it can be inferred that discrete tokenization is a very effective way to eliminate the dataset gap. Moreover, autoregressive models based on cross-entropy classification loss are key to the scalability of trajectory generation models, which aligns with why large language models (LLMs) have significant scaling capabilities.
@@
Table 7: Comparison of different tokenizer. Experimental results are based on the SMART 7M
@@
SMART 1M 0.423 0.782 0.835 8hours 10.30ms
SMART 7M 0.436 0.809 0.852 23hours 17.21ms
SMART 26M 0.442 0.817 0.864 3days 25.94ms
SMART 101M 0.457 0.819 0.872 1week 46.58ms
The results in Table 9 highlight the performance of SMART models with different parameter scales across various metrics. As the model scale increases from SMART 1M to SMART 101M, there is a significant improvement in both the interactive metrics and the map-based metrics. This indicates that larger models are better at capturing interactions and understanding map-based context, leading to enhanced performance in these areas. However, the kinematic metrics show minimal variation. Additionally, the training time and average inference time increase substantially with larger models, reflecting the trade-off between model performance and computational cost. Validation is conducted every 50,000 train steps. The model is considered to have converged if there is no significant loss reduction or metric improvement after five consecutive validations. The training and inference time is measured on 32 NVIDIA TESLA V100 GPUs.